OpenStudy (anonymous):

Solve 3c + 4 = -4c - 17 Answer 3 -3 21 -21

OpenStudy (anonymous):

What is it that you're stuck on? Do you know how to start?

OpenStudy (anonymous):

3c+4c=-17-4 combine like terms 7c=-21 divide by 7 c=-3

OpenStudy (anonymous):

3c + 4 = -4c - 17 3c + 4c + 4 = -4c + 4c - 17 7c + 4 = -17 7c + 4 - 4 = -17 - 4 7c = -21 7c ÷ 7 = -21 ÷ 7 c = -3

OpenStudy (anonymous):

tthanks and no i had no idea how to start it,

OpenStudy (anonymous):

WelcOme!

OpenStudy (anonymous):

np :)